Test the claim that the proportion of men who own cats is significantly different than the proportion of women who own cats at the 0.01 significance level.

The null and alternative hypothesis would be:

H0:pM=pFH0:pM=pF
H1:pM<pFH1:pM<pF

H0:μM=μFH0:μM=μF
H1:μM>μFH1:μM>μF

H0:pM=pFH0:pM=pF
H1:pM>pFH1:pM>pF

H0:pM=pFH0:pM=pF
H1:pM≠pFH1:pM≠pF

H0:μM=μFH0:μM=μF
H1:μM≠μFH1:μM≠μF

H0:μM=μFH0:μM=μF
H1:μM<μFH1:μM<μF



The test is:

two-tailed

left-tailed

right-tailed



Based on a sample of 20 men, 45% owned cats
Based on a sample of 40 women, 60% owned cats

The test statistic is: (to 2 decimals)

The p-value is: (to 2 decimals)

Based on this we:

  • Fail to reject the null hypothesis
  • Reject the null hypothesis
Answer #1

a)

H0 : pM=pF
H1 : pM≠pF

The test is two tailed

test statistics:

pcap = (p1 *n1+ p2*n2)/(n1+n2

= ( 0.45 * 20+ 0.60 * 40)/(20+40)
= 0.55

z = ( p1- p2)/sqrt(pcap *(1-pcap) *(1/n1+1/n2))
= ( 0.45 - 0.60)/sqrt(0.55*(1-0.55)/*(1/20+1/40))
= -1.10


The p value is 0.27

Based on this we:

Fail to reject the null hypothesis as p value . 0.01 significan celevel